Grade 12Physical Chemistry

(i) Calculate the number of electrons which will together weigh one gram. (ii) Calculate the mass and charge of one mole of electrons.

sudhanshu
12 Years agoGrade 12

2 Answers

Gaurav
11 Years ago
Hello Student
(i) Mass of one electron = 9.10939 × 10–31 kg
Number of electrons that weigh 9.10939 × 10–31 kg = 1
Number of electrons that will weigh 1 g = (1 × 10–3 kg) = 0.1098 × 10–3 + 31 = 0.1098 × 1028 = 1.098 × 1027
(ii) Mass of one electron = 9.10939 × 10–31 kg Mass of one mole of electron = (6.022 × 1023) × (9.10939 ×10–31 kg) = 5.48 × 10–7 kg
Charge on one electron = 1.6022 × 10–19 coulomb
Charge on one mole of electron = (1.6022 × 10–19 C) (6.022 × 1023) = 9.65 × 104 C
Kushagra Madhukar
5 Years ago
Dear student,
Please find the solution to your problem below.
 
(i) Mass of one electron = 9.10939 × 10–31 kg
Number of electrons that weigh 9.10939 × 10–31 kg = 1
Number of electrons that will weigh 1 g = (1 × 10–3 kg) = 0.1098 × 10–3 + 31 = 0.1098 × 1028 = 1.098 × 1027
(ii) Mass of one electron = 9.10939 × 10–31 kg
Mass of one mole of electron = (6.022 × 1023) × (9.10939 ×10–31 kg) = 5.48 × 10–7 kg
Charge on one electron = 1.6022 × 10–19 coulomb
Charge on one mole of electron = (1.6022 × 10–19 C) (6.022 × 1023) = 9.65 × 104 C
 
Hope it helps.
Regards,
Kushagra